What the fluid equivalents for mass and energy? (one of the answers is fluid density)
Answer #1

taking the analogy from Bernoulli's equation

P1 + 0.5pv1^2 = P2 + 0.5pv2^2 -------(1)

mgh1 + 0.5mv1^2 = mgh2 + 0.5mv2^2 -------(2)

fluid equivalent of mass is fluid density

fluid equivalent of energy is Pressure